OpenStudy (anonymous):

Solve the quadratic equation x2 - 4x + 57 = -5 by completing the square. A. x equals one plus or minus i times the square root of fifty B. x equals two plus or minus the square root of fifty eight C. x equals two plus or minus i times the square root of fifty eight D. x equals two plus or minus the square root of sixty two

OpenStudy (kirbykirby):

\( \large x^2-4x+57 = -5\\ \large x^2-4x+62 = 0\) complete the square by adding and subtracting \(\displaystyle \left(\frac{b}{2a}\right)^2=\left( \frac{-4}{2(1)}\right)^2=4\) \[\large \underbrace{x^2-4x+4}-4+62=0\\ \large \underbrace{(x-2)^2}+58=0 \] Consider this as a difference of squares (i.e. \(a^2-b^2 =(a-b)(a+b)\) ) by re-writing it as: \[\large (x-2)^2-(-58)=0\\ \large [(x-2)-\sqrt{-58}][(x-2)+\sqrt{-58}]=0\\ \large (x-2-i\sqrt{58})(x-2+i\sqrt{58})=0\] Hence: \(\large x=2+i\sqrt{58}\) \(\large x = 2 - i\sqrt{58}\)